OpenStudy (anonymous):

if g(x)=log6 (x^2) find g'(x)

OpenStudy (agent47):

Hint: log6 is a constant, use the power rule. This is really simple

OpenStudy (agent47):

what's the derivative of: y = 5*x^2

OpenStudy (anonymous):

\[is\ \it\ \log_6(x^2)\ or\ \log_{e\ or\ 10}6 * x^2 \]

OpenStudy (anonymous):

is it*

OpenStudy (anonymous):

sorry I lost internet for a bit:(

OpenStudy (anonymous):

Is this the problem? $$\log(6) \cdot x^2$$

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

To take the derivative of an exponential like $x^2$, we simply take the exponent to the coefficient and subtract one from the exponent. Ignore leading constants for now, look at this: